OpenStudy (anonymous):

Which number line represents all of the values of x for the equation x2 = 36? @mathstudent55 I will draw the number lines for you :)

OpenStudy (mathstudent55):

\(x^2 = 36\) Subtract 36 from both sides: \(x^2 - 36 = 0\) Factor the left side: \((x + 6)(x - 6) = 36\) Set each factor equal to zero and solve: \(x + 6 = 0 \) or \(x - 6 = 0\) \(x = -6\) or \(x = 6\)

OpenStudy (mathstudent55):

The answer is a number line with a closed dot on -6 and a closed dot on 6.

OpenStudy (anonymous):

ok thanks so much! :)